3x-4y=-15
5x+y=-2

Answer by mananth(16946)

3 x -4 y = -15 .............1
Total value
5 x + 1 y = -2 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 4
3 x -4 y = -15
20 x + 4 y = -8
Add the two equations
23 x = -23
/ 23
x = -1
plug value of x in (1)
3 x -4 y = -15
-3 -4 y = -15
-4 y = -15 + 3
-4 y = -12
y = 3